# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v1

@@ 6 Concluding remarks
  • In the 1D Heisenberg $J_{1}J_{2}J3$ setting with $N=30$ spins, $J_{1}=1.0$ (fixed) and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2)$, $(0.2,0.5)$, $(0.5,0.2)$, hyperbolic GRU definitively outperforms Euclidean GRU four of four times. In this setting, the $J_{1}J_{2}J_{3}$ Hamiltonian exhibits even more pronounced hierarchical interaction structures comprising the first, second and third nearest neighbors. This is our third clue.
Given these results, and the established results in natural language processing that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical, tree-like structures [30], [35], [33], it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. More works are needed to confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
@@ Appendix A Convergence curves
Figure 11: Convergence curves of the energy variances using eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 in VMC runs for 1D TFIM with various $N$ (clockwise from top: $N=20,40,100,80$ spins). Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
@@
Figure 13: Convergence curves of the energy variances using 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with ($N_{x},N_{y}$) = (5,5), (7,7), (9,9), (8,8) lattices (clockwise from top). Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
@@
Figure 15: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
@@
Figure 16: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (5,5), (7,7), (9,9), (8,8) - for $N=30$ spins.
Figure 17: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D $J_{1}J_{2}J_{3}$ Heisenberg model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (5,5), (7,7), (9,9), (8,8) - for $N=30$ spins.

### v2

@@ 6 Concluding remarks
  • In the 1D Heisenberg $J_{1}J_{2}J_{3}$ setting with $N=30$ spins, $J_{1}=1.0$ (fixed) and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2)$, $(0.2,0.5)$, $(0.5,0.2)$, hyperbolic GRU definitively outperforms Euclidean GRU four out of four times in the first set of experiments (where the networks have the same number of parameters), and three out of four times in the second set of VMC experiments (where the hyperbolic GRU network has around 10-20 percent fewer parameters than the Euclidean GRU network). In this setting, the $J_{1}J_{2}J_{3}$ Hamiltonian exhibits even more pronounced hierarchical interaction structures comprising the first, second and third nearest neighbors. This is our third clue.
Given these findings, and the established results in natural language processing that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical, tree-like structures [30], [35], [33], it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. More works are needed to confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
@@ Appendix A Convergence curves
Figure 12: Convergence curves of the energy variances using eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 in VMC runs for 1D TFIM with various $N$ (clockwise from top: $N=20,40,100,80$ spins). The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.
@@
Figure 14: Convergence curves of the energy variances using 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with ($N_{x},N_{y}$) = (5,5), (7,7), (9,9), (8,8) lattices (clockwise from top). The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.
@@
Figure 16: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins. The red dashed line corresponds to the line of zero variance.
@@
Figure 17: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (0.0,0.5), (0.2,0.2), (0.5,0.2), (0.2,0.5) - for $N=30$ spins.
Figure 18: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D $J_{1}J_{2}J_{3}$ Heisenberg model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (0.0,0.5), (0.2,0.2), (0.5,0.2), (0.2,0.5)- for $N=30$ spins. The red dashed line corresponds to the line of zero variance.
Figure 19: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from top to bottom, ($J_{2},J_{3}$) = (0.2,0.5), (0.5,0.2) - for $N=30$ spins. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact energy line.
Figure 20: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D $J_{1}J_{2}J_{3}$ Heisenberg model with different $(J_{2},J_{3})$ values - from top to bottom, ($J_{2},J_{3}$) = (0.2,0.5), (0.5,0.2) - for $N=30$ spins. The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.